Express (7,0,9) as a linear combiantion of v1(1,1,1) v2(1,1,0) and v3(1,0,0)


1
Expert's answer
2021-08-04T14:57:17-0400

Let "u=(7, 0, 9), u=av_1+bv_2+cv_3." Then


"\begin{matrix}
 a(1)+b(1)+c(1)=7 \\
 a(1)+b(1)+c(0)=0 \\
 a(1)+b(0)+c(0)=9 \\
\end{matrix}"


"\begin{matrix}
 c=7 -(a+b)\\
 b=-a \\
 a=9 \\
\end{matrix}"


"\begin{matrix}
 c=7 \\
 b=-9 \\
 a=9 \\
\end{matrix}"




"u=(7, 0, 9)=9v_1-9v_2+7v_3"